PROBLEM to apply Newton's Root Approximation Method to.
Consider the functions g(x) = x3 ; h(x) = -3x+6
Apply Newton's Method to approximate x value intersection of these two graphs.
A. Please begin your first guess of x 1 = 1,ca lculate up to x3 value.

Answer by rothauserc(4718)

g(x) = h(x)
:
g(x) - h(x) = 0
:
1) x^3 +3x -6 = 0
:
We want to solve equation 1 for values of x using Newton's Method
:
Newton's Method says
:
x(n+1) = x(n) - f(x(n))/f'(x(n))
:
we begin with x(1) = 1
:
Note f'(x) = 3x^2 +3
:
x(1) = 1 -(-2)/6 = 1 +(1/3) = 4/3 = 1.3333
:
x(2) = 1.3333 -(0.3701/8.3331) = 1.2889
:
x(3) = 1.2889 -(0.0079/7.9838) = 1.2879
:
***********************************************************************
Turns out that 1.2879 is a good approximation for the intersection
:
usually we stop the iterations when the difference of two iterations is
less than 0.0001
***********************************************************************
: